OpenStudy (anonymous):

suppose you had the equation 5x * 4x = 20x what are the x and y intercepts?

OpenStudy (anonymous):

question doesn't really make sense

OpenStudy (anonymous):

\[20x^2=20x\] \[20x^2-20x=0\] \[20x(x-1)=0\] \[x=0, x=1\]